OpenStudy (anonymous):

Hello- Having trouble with this problem, How I start solving (2x+26)°? It's part of a parallel line problem, and I'm having a major brain fart. Any help is appreciated :) Thanks!

OpenStudy (anonymous):

If it's y = 2x + 26 it is in point slope form. Go to y = 26 then go to the right one space and up two repeat that until you have two points then draw you're line.

OpenStudy (anonymous):

It looks like this, 

OpenStudy (anonymous):

@jamesaa

OpenStudy (anonymous):

3x+33 = 2x + 26 x + 33 = 26 x = -7

OpenStudy (anonymous):

Then plugin x and find the two angles

OpenStudy (anonymous):

whoops I copied down the problem wrong

OpenStudy (anonymous):

Thank you soo much! So the answer would be- 3-7+33 -4+33 29... Do I subtract that from 180°?

OpenStudy (anonymous):

3x-33 = 2x + 26 x - 33 = 26 x = 59

OpenStudy (anonymous):

Both of the angles are 144 degrees.

OpenStudy (anonymous):

To find the angle adjacent subtract 144 from 180

OpenStudy (anonymous):

Makes more sense! Okay, thank you :)